# Forecasting Multivariate Urban Data via Decomposition and Spatio-Temporal Graph Analysis

### v1

@@ 5.3 Experiments
We now evaluate the performance of our model (labelled DST in the tables) and compare it with strong baselines in three different settings. In all experiments, the length of the look-back window is 336, and the learning rate is set to 0.0001. The length of the forecast horizon is denoted in the second column of tables 2, 3 and 4. For our model, we decided to extract only one seasonal component due to the short length of the input. For instance, when the length of the look-back window is 336 hours (equivalent to two weeks), it is logical to extract daily seasonal patterns since multiple daily cycles are present in this period while extracting weekly patterns would be less meaningful considering the limited number of weekly cycles.
@@
Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.402 0.322 0.421 0.324 0.441 0.334 0.481 0.382 0.532 0.429 0.745 0.589 4.51 0.62 192 0.422 0.333 0.439 0.340 0.447 0.344 0.483 0.376 0.535 0.431 0.813 0.615 3.87 2.06 336 0.427 0.341 0.447 0.350 0.467 0.358 0.488 0.382 0.542 0.431 0.804 0.623 4.47 2.57 720 0.423 0.347 0.469 0.375 0.472 0.379 0.452 0.342 0.549 0.432 0.827 0.644 9.80 7.46 Carbon 96 0.468 0.501 0.476 0.503 0.467 0.492 0.512 0.539 0.519 0.547 1.128 0.780 -0.21 -1.83 192 0.534 0.545 0.541 0.544 0.540 0.547 0.674 0.621 0.584 0.580 1.243 0.829 1.11 0.37 336 0.589 0.581 0.601 0.595 0.597 0.594 0.703 0.659 0.631 0.629 1.328 0.865 1.34 2.19 720 0.721 0.656 0.731 0.669 0.729 0.661 0.878 0.738 0.723 0.650 1.489 0.927 1.10 0.76 Electricity 96 0.131 0.228 0.149 0.252 0.136 0.232 0.190 0.301 0.189 0.298 1.528 0.927 3.68 1.72 192 0.141 0.241 0.160 0.264 0.167 0.269 0.202 0.313 0.204 0.311 1.521 0.927 11.87 8.71 336 0.153 0.260 0.170 0.279 0.169 0.275 0.210 0.318 0.231 0.336 1.530 0.934 9.47 5.45 720 0.171 0.282 0.204 0.314 0.200 0.298 0.239 0.350 0.246 0.361 1.535 0.942 14.5 5.37 Weather 96 0.178 0.261 0.169 0.251 0.175 0.259 0.364 0.406 0.319 0.329 0.258 0.254 -5.33 -3.98 192 0.234 0.315 0.229 0.300 0.225 0.299 0.436 0.439 0.322 0.354 0.339 0.294 -4.00 -7.14 336 0.381 0.443 0.321 0.403 0.352 0.412 0.470 0.469 0.401 0.391 0.421 0.343 -18.69 -29.15 720 0.549 0.536 0.414 0.465 0.421 0.481 0.467 0.447 0.433 0.444 0.465 0.394 -32.60 -36.04
Table 2: Experimental results on real data in multivariate input, multivariate output setting. The results were averaged over five runs to eliminate the effect of randomness. The IMP column displays the error reduction of our proposed model compared to the best baseline. The best results are highlighted in bold and the second-best results are in underlined.
Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.864 0.703 0.874 0.699 0.872 0.701 0.947 0.714 0.914 0.695 1.579 0.869 0.92 -0.57 192 0.903 0.719 0.926 0.728 0.931 0.732 1.008 0.744 0.973 0.739 1.725 0.926 2.48 1.24 336 0.916 0.722 0.954 0.738 0.955 0.745 1.057 0.792 0.969 0.744 1.799 0.948 3.98 2.17 720 0.945 0.735 0.986 0.746 1.015 0.783 1.063 0.814 1.009 0.751 1.877 0.977 4.16 1.47 Carbon 96 0.232 0.370 0.235 0.375 0.244 0.381 0.731 0.858 0.418 0.561 1.771 1.039 1.28 1.33 192 0.248 0.389 0.273 0.406 0.284 0.412 0.779 0.871 0.445 0.594 1.853 1.077 9.16 4.19 336 0.266 0.410 0.307 0.430 0.305 0.427 0.801 0.896 0.457 0.612 1.919 1.106 12.79 3.98 720 0.282 0.425 0.368 0.476 0.364 0.469 0.818 0.914 0.551 0.687 2.025 1.141 22.53 9.38 Electricity 96 0.148 0.273 0.166 0.292 0.169 0.293 0.194 0.309 0.200 0.291 1.221 0.858 10.84 6.19 192 0.176 0.298 0.189 0.312 0.191 0.320 0.211 0.319 0.209 0.319 1.223 0.857 6.87 4.49 336 0.188 0.310 0.209 0.329 0.204 0.321 0.218 0.330 0.211 0.334 1.297 0.881 7.84 3.43 720 0.214 0.335 0.252 0.371 0.255 0.374 0.240 0.351 0.258 0.359 1.369 0.908 10.83 4.56 Weather 96 0.096 0.206 0.127 0.234 0.131 0.237 0.109 0.212 0.112 0.243 0.130 0.262 11.93 2.83 192 0.378 0.416 0.412 0.425 0.423 0.427 0.401 0.419 0.439 0.461 0.451 0.478 5.74 0.72 336 0.780 0.604 0.834 0.613 0.845 0.621 0.806 0.605 0.812 0.604 0.844 0.620 3.23 0.00 720 1.237 0.797 1.408 0.825 1.392 0.818 1.398 0.816 1.418 0.876 1.470 0.934 11.14 2.33
@@
Table 3: Experimental results on real data in multivariate input, univariate output setting. The results were averaged over five runs to eliminate the effect of randomness. The IMP column displays the error reduction of our proposed model compared to the best baseline. The best results are highlighted in bold and the second-best results are in underlined.
Dataset Length DST DLinear NLinear Autoformer TimesFM Repeat Last IMP MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.867 0.695 0.874 0.699 0.872 0.701 0.961 0.732 0.917 0.699 1.579 0.869 0.57 0.57 192 0.921 0.724 0.926 0.728 0.931 0.732 1.023 0.760 0.971 0.741 1.725 0.926 0.54 0.55 336 0.950 0.736 0.954 0.738 0.955 0.745 1.084 0.803 1.001 0.747 1.799 0.948 0.42 0.27 720 0.984 0.747 0.986 0.746 1.015 0.783 1.092 0.833 1.013 0.755 1.877 0.977 0.20 -0.13 Carbon 96 0.216 0.357 0.235 0.375 0.244 0.381 0.615 0.796 0.422 0.567 1.771 1.039 8.08 4.80 192 0.264 0.398 0.273 0.406 0.284 0.412 0.631 0.818 0.446 0.601 1.853 1.077 3.29 1.97 336 0.293 0.422 0.307 0.430 0.305 0.427 0.722 0.846 0.463 0.615 1.919 1.106 3.3 1.17 720 0.360 0.471 0.368 0.476 0.364 0.469 0.779 0.875 0.557 0.690 2.025 1.141 1.10 -0.42 Electricity 96 0.154 0.277 0.166 0.292 0.169 0.293 0.198 0.311 0.209 0.291 1.221 0.858 7.23 4.81 192 0.173 0.297 0.189 0.312 0.191 0.320 0.214 0.323 0.223 0.342 1.223 0.857 8.46 4.81 336 0.196 0.315 0.209 0.329 0.204 0.321 0.231 0.348 0.231 0.345 1.297 0.881 3.92 1.87 720 0.238 0.357 0.252 0.371 0.255 0.374 0.288 0.389 0.307 0.398 1.369 0.908 5.55 3.77 Weather 96 0.119 0.228 0.127 0.234 0.131 0.237 0.138 0.245 0.157 0.282 0.130 0.262 6.30 2.56 192 0.421 0.433 0.412 0.425 0.423 0.427 0.476 0.501 0.470 0.483 0.451 0.478 -2.18 -1.88 336 0.843 0.621 0.834 0.613 0.845 0.621 0.864 0.658 0.872 0.684 0.844 0.620 -1.07 -1.30 720 1.414 0.826 1.408 0.825 1.392 0.818 1.412 0.835 1.555 0.954 1.470 0.934 -1.58 -0.97
Table 4: Experimental results on real data in univariate input, univariate output. The results were averaged over five runs to eliminate the effect of randomness. The IMP column displays the error reduction of our proposed model compared to the best baseline. The best results are highlighted in bold and the second-best results are in underlined.
As shown in Tables 2, 3, and 4, our proposed model has achieved the best or near-best results across all settings. We present detailed results in three different settings to ensure a fair comparison of the strengths and weaknesses of the baseline models and to clearly explain the effectiveness of our proposed model.
@@
DLinear and NLinear achieve the best performance among the baselines. Despite their impressive performance, they are univariate models that lack a mechanism to extract spatial dependencies. This limitation becomes evident in the multivariate input and univariate forecasting setting, where forecasting a variate may benefit from capturing lagged dependencies with other variates. Univariate models like DLinear or NLinear cannot leverage this information, leading to inferior performance compared to models that can, such as our proposed model and transformer-based models. This is particularly evident in the forecasting of PM2.5 with multivariate input.
@@
Transformer-based forecasting models, such as Autoformer, Fedformer, and Informer, are highly complex and over-parameterized, making them susceptible to overfitting. This issue is apparent in the forecasting of the carbon dataset. Additionally, we have included the inference results of TimesFM, which were obtained without additional training or fine-tuning. This lack of fine-tuning likely contributes to its suboptimal performance, as it might not have adapted well to the specific characteristics of the datasets. Pretrained models are generally not tailored to the nuances of new datasets unless they undergo further training to update their parameters. Consequently, the model’s inability to learn and adjust to the unique temporal patterns and data distribution of the datasets has likely resulted in less accurate forecasts compared to the other models that were specifically trained.
@@
Figure 8 presents a comparative analysis of the Mean Squared Error (MSE) performance of our proposed model, DST, against several baseline models across four datasets: Beijing PM2.5, Carbon, Electricity, and Weather. This evaluation is conducted on a multivariate input-multivariate output forecasting task, where the input and output lengths are set to 336 and 96, respectively. The results demonstrate that DST consistently achieves lower MSE values compared to the baselines, indicating its superior forecasting capability. Notably, traditional transformer-based models, such as Autoformer and Informer, tend to exhibit higher variance and median errors, suggesting increased instability in their predictions. DLinear, a linear baseline, performs competitively but is outperformed by DST, particularly on the Carbon dataset. The y-axis scales are adjusted for better visualization of performance differences among the models
6 Ablations
To thoroughly examine the properties of our proposed model, we conducted ablation studies to evaluate the effectiveness of each of its key modules. The modules under evaluation include (1) Decomposition, (2) Date and Time Embedding, (3) Temporal Feature Extraction, and (4) Spatial Feature Extraction. For each dataset, we compare the model’s performance with and without each of these modules.
@@ 6 Ablations
Due to the modular and pipelined structure of the proposed architecture, we can easily remove individual modules without impacting the functionality of the remaining components. The only exception is the Date and Time Embedding module, whose output is concatenated with the extracted spatio-temporal features. To simulate the removal of this module, we replace its output with zero-vectors, thereby mitigating its effect on the model’s performance.
@@
Table 5 demonstrates the effectiveness of each module within the model architecture. Among all the components, the temporal feature extraction module plays the most influential role, as indicated by the highest increase in Mean Squared Error (MSE) when this module is removed. This suggests that capturing temporal dependencies is crucial for accurately modeling the datasets, particularly for data characterized by strong temporal patterns, such as air pollution, carbon emissions, and electricity usage. The ablation results also reveal that other modules, such as decomposition, date and time embedding, and spatial feature engineering, contribute significantly to model performance, albeit to a lesser extent. These components work synergistically to enhance predictive accuracy by providing complementary information that captures different aspects of the data, such as spatial correlations and temporal variations. Overall, the study reveals the necessity of taking a multi-faceted approach that integrates various techniques to effectively handle complex, real-world datasets.
@@
Additionally, we conducted experiments to evaluate the impact of the number of GATv2 layers. Across all experiments, a single layer for spatial feature extraction consistently yielded the best performance. Figure 9 illustrates how the MSE changes as more GATv2 layers are added for the carbon intensity dataset.
@@
w/o Decomposition 0.441 0.474 0.154 0.179
w/o Date&Time Emb. 0.424 0.482 0.149 0.158
w/o Temporal F.E. 0.452 0.491 0.162 0.161
w/o Spatial F.E. 0.439 0.488 0.156 0.165
Original 0.404 0.460 0.133 0.155
@@ 7 Conclusion
In this paper, we introduced a novel multivariate time-series forecasting model that leverages the advanced capabilities of GATv2 to capture spatial dependencies among different time-series variables. Our approach, which incorporates a decomposition-based preprocessing step, has demonstrated incremental improvements in forecast accuracy and reliability across various real-world datasets. The experimental results underscore the potential of our model to outperform existing state-of-the-art methods, particularly in complex domains such as energy management, weather prediction, carbon intensity forecasting, and air pollution monitoring. By effectively modeling the intricate interdependencies within multivariate time-series data, our proposed model provides a robust framework for long-term forecasting, contributing to the development of smart infrastructure systems and enhancing societal outcomes. Future work will focus on further optimizing the model, exploring its applicability to additional domains, and integrating adaptive graph learning techniques to enhance its performance in dynamic environments.
Appendix
Appendix A Number of GATv2 layers
Figure 9: Changes of MSE loss by increasing the number of GATv2 layers used for spatial feature extraction on carbon intensity dataset.

### v2

@@ 5 Results
We now evaluate the performance of our model (labelled DST in Table 2) and compare it with strong baselines in three different settings. In all experiments, the length of the look-back window is 336, and the learning rate is set to 0.0001. The length of the forecast horizon is specified in the second column of Table 2. For our model, we decided to extract only one seasonal component due to the short length of the input. For instance, when the length of the look-back window is 336 hours (equivalent to two weeks), it is logical to extract daily seasonal patterns since multiple daily cycles are present in this period while extracting weekly patterns would be less meaningful considering the limited number of weekly cycles.
We consider 4 long-term forecast horizons, namely 96, 192, 336, and 720, as they are commonly used in long-term time-series forecasting studies Wu et al. (2021); Zhou et al. (2022); Zeng et al. (2023). These multi-day to multi-week horizons are also more suitable for downstream applications, such as generator scheduling, carbon-aware computing Maji et al. (2022), and city-scale planning Mouatadid et al. (2024), where long-range forecasts support critical decision-making tasks more effectively than short-term ones.
@@
Dataset Look-ahead (# samples) DST s-Mamba DLinear Autoformer TimesFM PatchTST Repeat Last MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE Air pollution 96 0.864 0.703 0.874 0.699 0.872 0.701 0.947 0.714 0.914 0.695 0.901 0.699 1.579 0.869 192 0.903 0.719 0.926 0.728 0.931 0.732 1.008 0.744 0.973 0.739 0.944 0.735 1.725 0.926 336 0.916 0.722 0.954 0.738 0.955 0.745 1.057 0.792 0.969 0.744 0.958 0.753 1.799 0.948 720 0.945 0.735 0.986 0.746 1.015 0.783 1.063 0.814 1.009 0.751 0.989 0.803 1.877 0.977 Carbon 96 0.232 0.370 0.235 0.375 0.244 0.381 0.386 0.506 0.418 0.561 0.349 0.497 1.771 1.039 192 0.248 0.389 0.273 0.406 0.284 0.412 0.371 0.491 0.445 0.594 0.413 0.569 1.853 1.077 336 0.266 0.410 0.307 0.430 0.305 0.427 0.454 0.532 0.457 0.612 0.435 0.588 1.919 1.106 720 0.282 0.425 0.368 0.476 0.364 0.469 0.407 0.502 0.551 0.687 0.512 0.617 2.025 1.141 Electricity 96 0.148 0.273 0.166 0.292 0.169 0.293 0.194 0.309 0.200 0.291 0.198 0.311 1.221 0.858 192 0.176 0.298 0.189 0.312 0.191 0.320 0.211 0.319 0.209 0.319 0.214 0.326 1.223 0.857 336 0.188 0.310 0.209 0.329 0.204 0.321 0.218 0.330 0.211 0.334 0.237 0.349 1.297 0.881 720 0.214 0.335 0.252 0.371 0.255 0.374 0.240 0.351 0.258 0.359 0.284 0.397 1.369 0.908 Weather 96 0.096 0.206 0.127 0.234 0.131 0.237 0.109 0.212 0.112 0.243 0.115 0.256 0.130 0.262 192 0.378 0.416 0.412 0.425 0.423 0.427 0.401 0.419 0.439 0.461 0.437 0.459 0.451 0.478 336 0.780 0.604 0.834 0.613 0.845 0.621 0.806 0.605 0.812 0.605 0.829 0.618 0.844 0.620 720 1.237 0.797 1.408 0.825 1.392 0.818 1.398 0.816 1.418 0.876 1.378 0.133 1.470 0.934
@@
Table 2: Comparison of the forecasting accuracy of DST and baselines on four dataset in the multivariate input, univariate output setting. The results were averaged over five runs to mitigate the effect of randomness. The best results are highlighted in bold and the second-best results are underlined.
As shown in Table 2, DST achieves the best performance, with respect to MSE and MAE, in almost all settings. Notably, it outperforms the best baseline model, which is s-Mamba in most cases, by up to 9.10% on average. It can be seen that the gap between DST and the second best method generally widens for longer forecast horizons. This highlights the strength of the proposed framework for long-term multivariate time-series forecasting.
@@
To further demonstrate the effectiveness of DST, we provide a visual comparison of its predictions against those of the best-performing baselines, s-Mamba and DLinear, in Figure 7. These examples highlight the ability of our model to more accurately capture the temporal dynamics of the target time-series across different domains. In particular, DST demonstrates smoother forecasts and better alignment with the ground truth values. These qualitative results complement the quantitative improvements reported in Table 2, offering additional insight into the predictive strengths of our approach.
(a) Air Pollution
(b) Carbon Intensity
(c) Electricity
(d) Weather
Figure 7: Comparison between the forecasts of DST, and two best performing baselines: s-Mamba and DLinear, on selected samples from a randomly chosen variate in each dataset.
@@
Apart from s-Mamba, DLinear also exhibits reasonable performance among the baselines. Despite its good performance, it is a univariate model that lacks a mechanism to extract spatial dependencies. This limitation becomes evident in the multivariate input and univariate forecasting setting, where forecasting a variate may benefit from capturing lagged dependencies with other variates. Univariate models cannot leverage this information, leading to inferior performance compared to the models that can, such as our proposed model (DST) and transformer-based models. This is particularly evident in the forecasting of PM2.5 air pollution dataset.
@@
Transformer-based forecasting models, such as Autoformer, Fedformer, and Informer, are highly complex and over-parameterized, making them susceptible to overfitting. This issue is apparent in the forecasting of carbon intensity. Additionally, we have included the forecasting performance of TimesFM, which was obtained without fine-tuning (i.e. its zero-shot performance, which was also reported in Das et al. (2024)). This lack of fine-tuning likely contributes to its subpar performance, as it might not have adapted well to the specific characteristics of each dataset. Pretrained models must typically undergo additional training to align with the unique temporal patterns and distributions of new datasets. Without this adaptation, performance may suffer in comparison to models trained directly on the target datasets.
@@
Figure 6 presents a comparative analysis of the forecasting performance, in terms of the distribution of MSE, of our proposed model, DST, against several baseline models across four datasets: Air pollution, Carbon, Electricity, and Weather. The observed variability in MSE across runs primarily stems from stochastic elements in the training process, including random weight initialization, and data shuffling due to mini-batch sampling. One baseline model, TimesFM, is excluded from the plots because it is evaluated solely in the zero-shot inference setting. Consequently, it produces the MSE value that is reported in Table 2, with no variability. This evaluation is conducted on a multivariate input-multivariate output forecasting task, where the input and output lengths are set to 336 and 96, respectively. The results demonstrate that DST consistently achieves lower MSE values compared to the baselines, indicating its superior forecasting capability. Transformer-based models, such as Autoformer and PatchTST, exhibit relatively higher median errors and greater variability. DLinear and s-Mamba perform relatively well, but are still outperformed by DST in all datasets. We note that the y-axis is exaggerated in some cases to better highlight performance differences between models.
6 Ablation Study
@@
We conducted an ablation study to evaluate the effectiveness of each module in DST to better explain its performance. The modules under evaluation are (1) Decomposition, (2) Date and Time Embedding, (3) Temporal Feature Extraction, and (4) Spatial Feature Extraction. For each dataset, we compare the model’s performance without each of these modules. Due to the modular structure of DST, we can easily remove individual modules without impacting the functionality of the remaining components. The only exception is the Date and Time Embedding module, whose output is concatenated with the extracted spatio-temporal features. To remove the influence of this module, we replace its output with a zero vector.
@@
Table 3 demonstrates the effectiveness of each module within the model architecture. Among all the components, the temporal feature extraction module plays the most influential role, as indicated by the highest increase in MSE when this module is removed. This suggests that capturing temporal dependencies is crucial for accurately modeling the datasets, particularly for data characterized by strong temporal patterns, such as air pollution, carbon emissions, and electricity usage. Decomposition is also among the most effective components, especially for air pollution and weather forecasting. The ablation results also reveal that other modules, i.e. date and time embedding and spatial feature extraction, contribute significantly to model performance, albeit to a lesser extent. These components work synergistically to enhance predictive accuracy by providing complementary information that captures different aspects of the data, such as spatial correlations and temporal variations. Overall, the study reveals the necessity of taking a multi-faceted approach that integrates various techniques to effectively handle complex, real-world datasets.
@@
Additionally, we carried out experiments to evaluate the impact of the number of GATv2 layers on DST’s performance. We found that across all experiments, a single layer for spatial feature extraction consistently yields the best performance. As a result, DST is implemented with a single GATv2 layer.
@@
DST w/o Decomposition 0.441 0.474 0.154 0.179
DST w/o Date&Time Emb. 0.424 0.482 0.149 0.158
DST w/o Temporal F.E. 0.452 0.491 0.162 0.161
DST w/o Spatial F.E. 0.439 0.488 0.156 0.165
DST 0.404 0.460 0.133 0.155
@@ 7 Conclusion
In this paper, we introduced a novel multivariate time-series forecasting model that leverages a GNN architecture integrating GATv2 and TCN. This GNN relies on a distinct graph structure per time-series component. To learn this graph structure, we implemented a preprocessing step to decompose the time-series into its fundamental components and proposed an effective algorithm for learning the per-component graph structure, based on historical sensor data. Extensive experiments on four real-world urban datasets with different temporal resolutions, numbers of sensors, and interdependencies confirm that DST consistently achieves the best long-term forecasting performance across different forecast horizons. Specifically, it outperforms the best-in-class time-series forecasting model by 2.89% to 9.10% in different datasets.
Future work will focus on using these forecasting models in various decision-making tasks in the urban environment, from infrastructure planning to assignment of AI workloads (particularly training and fine-tuning) to datacenters, to quantify improvements in decision quality resulting from more accurate long-term forecasts. Furthermore, we intend to explore how training the long-term forecasting models using the task loss would change the decision quality.